OpenStudy (bbb911):

2.94 divided by 0.01332

pooja195 (pooja195):

220.720720721

OpenStudy (anonymous):

use calculator!!

OpenStudy (bbb911):

i need to see work before i use t he calcultor!!

OpenStudy (bbb911):

what i am having trouble is were i ca put the decimal in 0.01332

OpenStudy (saifoo.khan):

The decimal is already there?

OpenStudy (bbb911):

okay you know how you have to move the decimals ? so on 2.96 i move decimal twice and it looks like this 294.0 right?

OpenStudy (bbb911):

i mean 296.0

OpenStudy (saifoo.khan):

YES! You're on the right track. you moved the decimal of numerator twice on the right, so you will also move the decimal of denominator twice on the right.

OpenStudy (bbb911):

yes i did that to i go 1.332 right?

OpenStudy (saifoo.khan):

Yes. Now you have to remove the decimal from the denominator, so you'll do teh same thing which you did above.

OpenStudy (bbb911):

like when you said remove it do you mean by removing the decimal completely ? or?

OpenStudy (saifoo.khan):

yes, completely.

OpenStudy (bbb911):

so will i end up with 1332 right?

OpenStudy (saifoo.khan):

yep. and in the numerator..?

OpenStudy (bbb911):

296.00

OpenStudy (saifoo.khan):

\[\frac{2.94}{0.01332 } \to \frac{294}{1.332}\to \frac{294000}{1332}\]

OpenStudy (bbb911):

but i dont really have to include al those 000 at the end right?

OpenStudy (saifoo.khan):

You have to, because when the second time you move the decimal in the bottom, you have to move the move the decimal in the numerator too

OpenStudy (bbb911):

so when i divide is it going to be like this 296000 divided by 1332 or the other way around ?

OpenStudy (saifoo.khan):

Yes^

OpenStudy (bbb911):

kk thank you so much saifoo :)

OpenStudy (saifoo.khan):

You're welcome. (: